OpenStudy (anonymous):

which fraction would you use to find 40% of 57 percent metally?

Directrix (directrix):

40% = 40/100 = 4/10 = 2/5. 2/5 --> I don't know that that is a mental type help for me.

OpenStudy (anonymous):

Good call, Directrix. 2/5(57)=114/5 approx 115/5=23

OpenStudy (anonymous):

So forty percent of 57 percent is 23 percent.

OpenStudy (anonymous):

Approximately.....